Begin again by assuming 70. mg of caffeine is dissolved in 4.0 mL of water. This time, however, the aqueous solution will only be extracted ONCE with 6.0 mL of methylene chloride. Calculate the total milligrams of caffeine that can be extracted from the aqueous layer using this volume of methylene chloride. Caffeine has a distribution coefficient of 4.6 (K), between methylene chloride (CH2Cl2) and water.
Answer #1

Let x mg of caffeine is extracted from the aqueous layer to organic layer. The amount of caffeine in aqueous layer is (70.-x)mg .

Kd = 4.6 = [A]org / [A]aq =( x mg/6.0ml)/((70.-x)mg/4.0ml)

x/(210-3x) = 2.3

x = 2.3×210/7.9 = 61.14mg

The mass of caffeine that is extracted by given volume of methylenechloride = 61 mg (Answer)